OpenStudy (anonymous):

need help with these pairs this site is so great ove been getting really nice help from nice people. the ordered pairs (0,0) (2,4) (4,8) and (6,12) represent a funtion what is the rule of this function? A. y=x+2 B. y=3x C. y=x+4 D. y+2x @Marie_Loves_You @Rhystic

OpenStudy (anonymous):

@josedavid

OpenStudy (anonymous):

A or C

OpenStudy (wwwhhhaaattt?):

@Marie_Loves_You Do not give direct answers please Its against CoC

OpenStudy (anonymous):

im just looking for help can try and figure the rest out. not looking for an exact answer @WWWHHHAAATTT?

OpenStudy (josedavid):

is a

OpenStudy (wwwhhhaaattt?):

What do you think it is?

OpenStudy (anonymous):

im kind of leaning more towards A but im not positive can you help me work through it ? @WWWHHHAAATTT?

OpenStudy (rhystic):

You need to just plug in the the coordinates for the equation. (x,y)

OpenStudy (wwwhhhaaattt?):

Ok, First. Give me an Explanation of why you think its A.

OpenStudy (anonymous):

i have no idea @WWWHHHAAATTT?

OpenStudy (wwwhhhaaattt?):

Can you try at least?

OpenStudy (rhystic):

If you use the first set of coordinates you would be able to cancel a couple of the answers because the pair is (0,0) which means x = 0 and y = 0 [ (x,y) (0,0) ] if you plug these values in for the equations you should see that two of the equations will not work.

OpenStudy (anonymous):

ok so B and D wouldnt work because they dont fit with the coordinates? @Rhystic

OpenStudy (rhystic):

First off is D y= 2x or y +2x = 0? secondly how are you plugging in the numbers? because it seems like you're just spouting stuff based off what marie said which is wrong by the way.

OpenStudy (rhystic):

if you utilized (0,0) then B and D would be the only ones that work ...

OpenStudy (anonymous):

sorry @Rhystic just trying to figure this out.

OpenStudy (rhystic):

I'll do 0,0 for you.

OpenStudy (rhystic):

A. y = x +2 -> 0 = 0 +2 2 is not = 0 B. y = 3x -> 0 = 3(0) this works. C. y = x +4 -> 0 = 0 +4 0 is not = 4 D. y = 2x - > 0 = 2(0) this works. You now have B & D left that might be correct. Take the next set of coordinates : (x,y) = (2,4) and plug them into B & D and see which one works.

OpenStudy (wwwhhhaaattt?):

Dayum @Rhystic you good Almost to good

OpenStudy (anonymous):

its D y+2x

TheSmartOne (thesmartone):

And do what @rhystic did for all the point and just find an equation for which all these points work properly :)

OpenStudy (rhystic):

correct ally good job.

OpenStudy (anonymous):

thanks for all your help!!!1

OpenStudy (wwwhhhaaattt?):

SO GET OWNED @josedavid

OpenStudy (wwwhhhaaattt?):

You just got SHREKED